i want to check my answer on this one please tell me if i did it right.
2L+2W=72
L=2W+3
My answer is L=25 W=11
thanx!
~Carina~

Answer by ankor@dixie-net.com(22740)

2L+2W=72
L=2W+3
My answer is L=25 W=11
:
You're right, but you can check any solution by substituting your answers in the
original equations.
:
In the 1st equation:
2(25) + 2(11) = 72
50 + 22 = 72, proving that your solutions are correct.
:
In the 2nd equation: L = 2W + 3
25 = 2(11) + 3
25 = 22 + 3 also proves your answers